# Differentiable master equation solver for quantum device characterisation

### v1

@@ I Introduction
The challenge of characterising quantum devices has inspired a myriad of approaches, with the goal of furthering the understanding and application of quantum technologies gebhart2023learning . The simplest approach to characterising aspects of a quantum system is phenomenological fitting, which can yield useful parameter values with minimal resources and carefully selected measurements. More complex methods such as quantum state tomography leonhardt1995quantum and quantum process tomography mohseni2008quantum facilitate full characterisation of the states prepared in a device or how they are changed by device operation, but often require large amounts of data as well as access to observables providing a full tomographic basis set. Several approaches have been developed to improve the efficiency of these tomographic techniques, or to extend their descriptive power, by having appropriate models incorporated into the characterisation of both closed cole2005identifying ; devitt2006scheme ; zhang2014quantum and open quantum systems jorgensen2019exploiting ; white2022non .
@@
The success of modern deep learning methods is underpinned by differentiable programming. This programming paradigm facilitates the automatic differentiation required for training deep neural networks with many parameters, where numerical and symbolic derivatives become severely impractical rumelhart1986learning ; baydin2018automatic . Beyond applications to deep learning, differentiable programming has recently been applied to physical models in the context of physics-aware machine learning craig2024bridging ; pachalieva2022physics ; wright2022deep , efficient computation of gradients with respect to complex physical models liao2019differentiable ; Hu2020DiffTaichi: , and optimisation of non-equilibrium steady states vargas2021fully . There have also been recent advances in using differentiable models for simulations of quantum systems with a focus on optimal control schafer2020differentiable ; coopmans2021protocol ; khait2022optimal ; coopmans2022optimal ; sauvage2022optimal .
@@
To demonstrate the power and broad applicability of our approach, we have designed two case studies with different levels of complexity. The first case study of transport through a single quantum dot with an orbital excited state is designed to highlight our approach’s capabilities in handling differentiable and non-differentiable parameters. The second case study of transport through a double quantum dot coupled to a phonon bath is designed to explore characterisation of the environment gullans2018probing ; hofmann2020phonon ; barr2023spectral . In this second case study, we showcase the ability of our algorithm to discriminate between and correctly select from a set of models featuring different phonon spectral densities relating to different device materials and geometries. To quantitatively assess the performance of our approach, these case studies utilise simulated data. A sketch of the differentiable fitting process, which is central to our approach, is shown in Figure 1(a) along with both case study scenarios in (b) and (c), respectively.
@@
Figure 1: (a) A schematic of the process of estimating parameters of an open quantum system using gradient descent. Here, $H_{\mathrm{S}}$, $H_{\mathrm{E}}$ are the system and environment Hamiltonians, respectively, and $H_{\mathrm{I}}$ is the system-environment interaction Hamiltonian. Expectation values of an observable $\langle\hat{O}\rangle$ are generated from a ground truth model, which would be experimental data in practical applications, and a differentiable model is used to fit these values by gradient descent of a loss function. (b) Representation of the first case study of a single quantum dot with an excited state coupled to left and right leads. Here, $\mu_{\mathrm{L}(\mathrm{R})}$ is the chemical potential of the left(right) lead, $\Gamma_{\mathrm{L}(\mathrm{R})}$ is the tunnel rate between the dot states and the left(right) lead, and $E_{0(1)}$ is the electrochemical potential of the dot in the ground(excited) state. (c) Representation of the second case study of a double quantum dot coupled to left and right leads, and a phonon bath. In each case study the observable used to inform parameter estimation is the steady state current from left to right lead. Here $E_{\mathrm{L}(\mathrm{R})}$ is the chemical potential of the dot with a charge on the left(right) dot, $t_{c}$ is the tunnel coupling between the dots, and $\gamma_{p}$ is the coupling strength to the phonon bath.
@@ II Differentiable Master Equation Solver
Numerically solving such a quantum master equation using differentiable programming facilitates evaluation of the gradient of an observable with respect to the master equation parameters. To compute steady states, we solve the set of linear equations resulting from Eq. (1) under the condition $\dot{\rho}(t)=0$, and for time evolution we use ODE solvers as discussed in Section VII. A differentiable model allows gradients to be obtained with similar computation time to forward evaluation of the model using automatic differentiation, without the need for inefficient finite-element methods. Our implementation can deal with arbitrary Liouvillian superoperators, and is thus applicable to a range of physical systems. Related work has demonstrated differentiable computation of steady state solutions using JAX jax2018github for quantum heat transfer models vargas2021fully . Our approach differs in implementation, design, and application; we use TensorFlow, provide more general construction of Lindblad master equations, and embed our solver in a data-driven device characterisation algorithm.
@@ IV Optimising Non-Differentiable Parameters
To deal with a mixture of differentiable and non-differentiable parameters, we design an optimisation procedure which takes advantage of the fast differentiable model. The idea is to find the non-differentiable parameters for which the remaining differentiable parameters provide the best fit. A Nelder-Mead optimisation routine nelder1965simplex finds appropriate values for the non-differentiable parameters, where the function being optimised is a fit of the model to ground truth data using the remaining differentiable parameters. This fit first involves a log-spaced grid search of the differentiable elements of $\bm{\theta}^{\mathrm{v}}$ where the evaluation of $f(\bm{\theta})$ with the lowest loss with respect to the ground truth data is selected as a starting point for a short gradient descent optimisation using the Adam optimiser. The evaluation of the function being optimised by the Nelder-Mead routine is simply the final loss after this gradient descent. Finally, using the optimal non-differentiable parameters, conducting a longer gradient descent optimisation of the differentiable parameters provides an estimate of all parameters in $\bm{\theta}$. A schematic of this optimisation procedure is shown in Appendix B.
@@ V Case Study: Single Quantum Dot with Orbital Excited State
Figure 3: Distributions of percentage errors for parameter estimation over a test set of 100 randomly generated current traces for a single quantum dot with an orbital excited state. Distributions displayed are: (a)-(c) optimal values of non-differentiable parameters $[\mu_{l},\mu_{r},\delta]$, (d)-(f) optimal values of differentiable parameters $[\Gamma_{\mathrm{L}},\Gamma_{\mathrm{R}},T]$, and (g)-(i) posterior means for the differentiable parameters. All distributions are peaked at zero error, with a small number of outliers. Insets show the target data (red points) and optimal fit (blue line) for selected outliers.
An example of our parameter estimation algorithm is shown in Figure 2 with parameter estimates shown in Table 1. In this example, the largest absolute percentage error in optimised non-differentiable values is 1.2%, and in optimal differentiable values is 4.5%. Posterior means and standard deviations drift slightly from the optimal parameter values, but retain a good fit to the data. In Figure 3 we consider the distributions of percentage errors for each parameter for a set of 100 iterations of our parameter estimation algorithm on randomly generated data. The distributions are all peaked at zero percentage error, indicating that our parameter estimation method is generally effective. There are a small number of outliers with large errors, but upon further inspection we find that most of these outliers have no discernible second peak in the current profile (see insets of Figure 3, specifically Outliers A, C, and D). This means the location of the excited state cannot be determined at the first optimisation step, which impacts the estimation of further parameters. The outliers have a lower signal to noise ratio than typical samples which may further impact the parameter estimation, as evident in Outlier B which does have an excited state peak. From Figure 3, we observe that parameter estimate using the posterior mean from HMC samples has a more narrow distribution around zero percentage error, and is therefore generally better than the gradient descent optimal values.
@@ VI Case Study: Double Quantum Dot
where $\omega_{d}=\frac{2\pi c}{d}$ and $\omega_{a}=\frac{2\pi c}{a}$, with $d$ the centre-to-centre separation of the dots, $a$ the linear extent of the dots, and $c$ the speed of sound in the crystal, and $J_{0}$ is a scale factor. We also have $n$ as the dimension of the system, and $\alpha=0$ for piezoelectric coupling and $\alpha=2$ for deformation potential coupling to phonons. The complete set of spectral densities for varying dimensions and coupling mechanism is shown in Appendix D.
@@
We now consider the model presented above to assess the performance of our approach on a more complex system. This model has a single controlled parameter $\bm{\theta}^{\mathrm{c}}=[\epsilon]$ which is varied such that each dot energy level is swept through the bias window. We choose the variable parameters to be $\bm{\theta}^{\mathrm{v}}=[\mu_{l},\mu_{r},\Gamma_{\mathrm{L}},\Gamma_{\mathrm{R}},t_{c},J_{0},T]$ which can be separated into differentiable parameters $[\Gamma_{\mathrm{L}},\Gamma_{\mathrm{R}},t_{c},J_{0},T]$ and non-differentiable parameters $[\epsilon_{0},\epsilon_{\mathrm{bias}}]$ which are the points where $\epsilon=0$ and $\epsilon=\mu_{\mathrm{L}}-\mu_{\mathrm{R}}$, respectively. The remaining parameters are set to representative fixed values of $c_{s}=3000\ \mathrm{ms}^{-1}$, $a=20\ \mathrm{nm}$, and $d=100\ \mathrm{nm}$. We choose these parameters to be fixed as they can all be estimated from the device design, but they could also be included in the set of differentiable parameters. The dot radius $a$ in particular has minimal impact on the fit for the energy scales probed by typical experimental measurements. The problem of estimating parameters for this model from a single current trace is expected to be underdetermined due to the number of parameters defining the system. We find fitting to data points is accurate, with an example shown in Figure 4. Discussion of parameter estimation for this model can be found in Appendix C.
Capitalising on the accuracy of fitting to data, we use our approach to characterise other aspects of a device hosting these dots. We focus on model selection for the interaction of the double quantum dot with phonons, using our fitting process to perform model selection on the phonon spectral density. The phonon spectral density is determined by material properties and dimensionality of the semiconductor in which the dots are defined, as well as the geometry of the dots, as outlined in Appendix D. Accurately distinguishing between piezoelectric and deformation potential phonon coupling would yield information about the device in which the dots are formed. Determining the dimensionality of the phonon spectral density would yield information about the confinement of phonon modes in the device and help identify any possible waveguide effects.
@@ VII Time Evolution
Having considered parameter estimation from steady-state solutions to quantum master equations using our differentiable solver, this section briefly explores time-evolution functionality and its limitations. Dynamic control of qubits is a natural motivation for studying time evolution, and thus we consider the model of a double quantum dot coupled to fermionic leads and a phonon bath outlined above as charge qubits have been implemented in such systems gorman2005charge ; petersson2010quantum . A Lindblad master equation can be recast as a set of coupled ordinary differential equations (ODEs) for time evolution solutions, which means standard ODE solvers can be used. Our solver implements a $4^{\mathrm{th}}$ order Runge-Kutta method butcher2016numerical to compute time-dynamics. To verify that our solver is accurate, Figure 7 shows that the steady state solution is reached after a suitable evolution time under the influence of a time-independent Hamiltonian. This figure also demonstrates batch computation of time-evolution, whereby a batch of $n_{\mathrm{b}}=250$ detuning values is evolved from an initial state of $|\psi(0)\rangle=(|0\rangle+|L\rangle+|R\rangle)/\sqrt{3}$ over 15 ns for 1500 time steps in 0.38s using the same hardware as discussed in Appendix A.
@@ VIII Conclusion
We have presented a fast and differentiable quantum master equation solver, and demonstrated its capability to compute steady state solutions and time evolution of open quantum systems. Our solver can be applied to any system described by a Lindblad master equation, and minimal changes would permit application to a more general time-local approach involving fewer approximations. The speed of our solver comes from its capability to compute a batch of outputs in a single evaluation of a model. This speed in turn facilitates more involved algorithms requiring many evaluations, which would be impractical using standard solvers. An example of such an algorithm is the model fitting approach we develop which employs both gradient-based and gradient-free optimisation, and MCMC methods.
Our parameter estimation algorithm has been applied to case studies of transport through a single quantum dot with an orbital excited state, and a double quantum dot coupled to a phonon bath. We have found our approach to be successful for dealing with parameter estimation in the case of a single quantum dot. To highlight the utility of our algorithm, we have also performed model selection to determine the electron-phonon coupling mechanism in a double quantum dot. Our approach is successful in reliably determining the ground truth spectral density. Our approach could be directly applied to experimental studies of the spectral density, such as in Ref. hofmann2020phonon . Decoherence from electron-phonon interactions impacts both charge and spin qubits, therefore characterising the details of this coupling mechanism is important for understanding qubit dynamics.
Our differentiable master equation solver and its application to parameter estimation is a new approach to characterising quantum systems, and we anticipate similar differentiable models will provide significant utility across a range of applications. Physics-aware machine learning is a notable area in which differentiable models have been applied previously craig2024bridging ; pachalieva2022physics ; wright2022deep , and our solver stands to widen the impact of such approaches. In addition, differentiable pulse simulators have been identified as ‘urgently needed’ in the field of quantum pulse engineering liang2022variational . Our work steps towards reaching this goal in the framework of differentiable programming for quantum master equations, but there remains scope for future work for fully differentiable time evolution.

### v2

@@ I Introduction
The challenge of characterising quantum devices has inspired a myriad of approaches, with the goal of furthering the understanding and application of quantum technologies gebhart2023learning . The simplest approach to characterising aspects of a quantum system is phenomenological fitting, which can yield useful parameter values with minimal resources and carefully selected measurements. Phenomenological fitting has the advantage of being readily applied to suitable data, but neglects detailed structure of the model describing a physical system and often fails when closed-form analytic solutions are unknown. More complex methods such as quantum state tomography leonhardt1995quantum and quantum process tomography mohseni2008quantum facilitate full characterisation of the states prepared in a device or how they are changed by device operation, but often require large amounts of data as well as access to observables providing a full tomographic basis set. Several approaches have been developed to improve the efficiency of these tomographic techniques, or to extend their descriptive power, by having appropriate models incorporated into the characterisation of both closed cole2005identifying ; devitt2006scheme ; zhang2014quantum and open quantum systems jorgensen2019exploiting ; white2022non .
@@
The success of modern deep learning methods is underpinned by differentiable programming. This programming paradigm facilitates the automatic differentiation required for training deep neural networks with many parameters, where numerical and symbolic derivatives become severely impractical rumelhart1986learning ; baydin2018automatic . Beyond applications to deep learning, differentiable programming has recently been applied to physical models in the context of physics-aware machine learning craig2024bridging ; pachalieva2022physics ; wright2022deep , efficient computation of gradients with respect to complex physical models liao2019differentiable ; Hu2020DiffTaichi: , optimisation of non-equilibrium steady states vargas2021fully , and characterisation of non-Markovian dynamics krastanov2020unboxing . Deep learning methods have recently been combined with differentiable models of quantum systems to enhance quantum device characterisation genois2021quantum ; youssry2024experimental . There have also been recent advances in using differentiable models for simulations of quantum systems with a focus on optimal control schafer2020differentiable ; coopmans2021protocol ; wittler2021integrated ; khait2022optimal ; coopmans2022optimal ; sauvage2022optimal , a requirement for the future of pulse engineering liang2022variational .
Figure 1: (a) A schematic of the process of estimating parameters of an open quantum system using gradient descent. Here, $H_{\mathrm{S}}$, $H_{\mathrm{E}}$ are the system and environment Hamiltonians, respectively, and $H_{\mathrm{I}}$ is the system-environment interaction Hamiltonian. Expectation values of an observable $\langle\hat{O}\rangle$ are generated from a ground truth model, which would be experimental data in practical applications, and a differentiable model is used to fit these values by gradient descent of a loss function. (b) Representation of the first case study of a single quantum dot with an excited state coupled to left and right leads. Here, $\mu_{\mathrm{L}(\mathrm{R})}$ is the chemical potential of the left(right) lead, $\Gamma_{\mathrm{L}(\mathrm{R})}$ is the tunnel rate between the dot states and the left(right) lead, and $E_{0(1)}$ is the electrochemical potential of the dot in the ground(excited) state. (c) Representation of the second case study of a double quantum dot coupled to left and right leads, and a phonon bath. In each case study the observable used to inform parameter estimation is the steady state current from left to right lead. Here $E_{\mathrm{L}(\mathrm{R})}$ is the chemical potential of the dot with a charge on the left(right) dot, $t_{c}$ is the tunnel coupling between the dots, and $\gamma_{p}$ is the coupling strength to the phonon bath. (d) An illustrative schematic of our approach for optimising differentiable and non-differentiable parameters when fitting to data. Here, we denote non-differentiable parameters as $\bm{\theta}^{\mathrm{v}}_{\mathrm{nd}}$, differentiable parameters as $\bm{\theta}^{\mathrm{v}}_{\mathrm{diff}}$, and optimised parameters are indicated with an asterisk.
@@
To demonstrate the power and broad applicability of our approach, we have designed two case studies with different levels of complexity. The features probed in our case studies are not suitable for phenomenological fitting, and each case study is suitably complex that typical analytic solutions for quantum dots are not applicable. The first case study of transport through a single quantum dot with an orbital excited state is designed to highlight our approach’s capabilities in handling differentiable and non-differentiable parameters. The second case study of transport through a double quantum dot coupled to a phonon bath is designed to explore characterisation of the environment gullans2018probing ; hofmann2020phonon ; barr2023spectral . In this second case study, we showcase the ability of our algorithm to discriminate between and correctly select from a set of models featuring different phonon spectral densities relating to different device materials and geometries. To quantitatively assess the performance of our approach, these case studies utilise simulated data. A sketch of the differentiable fitting process, which is central to our approach, is shown in Figure 1(a) along with both case study scenarios in (b) and (c), respectively, and an outline of our approach in (d).
@@ II Differentiable Master Equation Solver
Numerically solving such a quantum master equation using differentiable programming facilitates evaluation of the gradient of an observable with respect to the master equation parameters. To compute steady states, we solve the set of linear equations resulting from Eq. (1) under the condition $\dot{\rho}(t)=0$, and for time evolution we use ODE solvers as discussed in Section VII. A differentiable model allows gradients to be obtained with similar computation time to forward evaluation of the model using automatic differentiation, without the need for inefficient finite-element methods. Solving the steady state directly from the Liouvillian is critical to the speed of our solver by avoiding the many operations and storing of gradients required for a long-time integration of an ODE. Our implementation can deal with arbitrary Liouvillian superoperators, and is thus applicable to a range of physical systems.
@@
Related works have explored differentiable master equation solvers for time dynamics krastanov2020unboxing , and incorporated differentiable ODE solvers for learning dynamics of quantum systems with recurrent neural networks genois2021quantum . The effectiveness of differentiable models of quantum systems enhanced by machine learning is further demonstrated in youssry2024experimental in the context of photonic devices. Our approach takes a different direction to these works by considering steady state solutions, and implementing a data efficient characterisation algorithm which also handles non-differentiable parameters. Other related work has demonstrated differentiable computation of steady state solutions using JAX jax2018github for quantum heat transfer models vargas2021fully .
@@ IV Optimising Non-Differentiable Parameters
To deal with a mixture of differentiable and non-differentiable parameters, we design an optimisation procedure which takes advantage of the fast differentiable model. The idea is to find the non-differentiable parameters for which the remaining differentiable parameters provide the best fit. A Nelder-Mead optimisation routine nelder1965simplex finds appropriate values for the non-differentiable parameters, where the function being optimised is a fit of the model to ground truth data using the remaining differentiable parameters. This fit first involves a log-spaced grid search of the differentiable elements of $\bm{\theta}^{\mathrm{v}}$ where the evaluation of $f(\bm{\theta})$ with the lowest loss with respect to the ground truth data is selected as a starting point for a short gradient descent optimisation using the Adam optimiser. The evaluation of the function being optimised by the Nelder-Mead routine is simply the final loss after this gradient descent. Finally, using the optimal non-differentiable parameters, conducting a longer gradient descent optimisation of the differentiable parameters provides an estimate of all parameters in $\bm{\theta}$. A schematic of this optimisation procedure is shown in Figure 1(d), and discussion of its scaling can be found in Appendix B.
@@ V Case Study: Single Quantum Dot with Orbital Excited State
Figure 3: Distributions of percentage errors for parameter estimation over a test set of 100 randomly generated current traces for a single quantum dot with an orbital excited state. Distributions displayed are: (a)-(c) optimal values of non-differentiable parameters $[\mu_{l},\mu_{r},\delta]$, (d)-(f) optimal values of differentiable parameters $[\Gamma_{\mathrm{L}},\Gamma_{\mathrm{R}},T]$, and (g)-(i) posterior means for the differentiable parameters. All distributions are peaked at zero error, with a small number of outliers. Insets show the target data (red points) and optimal fit (blue line) for selected outliers. Summary statistics of these distributions can be found in Appendix C.
An example of our parameter estimation algorithm is shown in Figure 2 with parameter estimates shown in Table 1. In this example, the largest absolute percentage error in optimised non-differentiable values is 1.2%, and in optimal differentiable values is 4.5%. Posterior means and standard deviations drift slightly from the optimal parameter values, but retain a good fit to the data. In Figure 3 we consider the distributions of percentage errors for each parameter for a set of 100 iterations of our parameter estimation algorithm on randomly generated data. The distributions are all peaked at zero percentage error, indicating that our parameter estimation method is generally effective. There are a small number of outliers with large errors, but upon further inspection we find that most of these outliers have no discernible second peak in the current profile (see insets of Figure 3, specifically Outliers A, C, and D). This means the location of the excited state cannot be determined at the first optimisation step, which impacts the estimation of further parameters. The outliers have a lower signal to noise ratio than typical samples which may further impact the parameter estimation, as evident in Outlier B which does have an excited state peak.
From Figure 3, we observe that parameter estimates using the posterior mean from HMC samples has a more narrow distribution around zero percentage error, and is therefore generally better than the gradient descent optimal values. The additional information contained in the posterior distribution such as spread relative to the mean, multi-modal structures, and correlations between parameters provide further insight into results of our approach. In Appendix C, we use the additional information provided by posterior distributions to filter results to remove uncertain estimates and ensure a higher accuracy in accepted results.
@@ VI Case Study: Double Quantum Dot
where $\omega_{d}=\frac{2\pi c}{d}$ and $\omega_{a}=\frac{2\pi c}{a}$, with $d$ the centre-to-centre separation of the dots, $a$ the linear extent of the dots, and $c$ the speed of sound in the crystal, and $J_{0}$ is a scale factor. We also have $n$ as the dimension of the system, and $\alpha=0$ for piezoelectric coupling and $\alpha=2$ for deformation potential coupling to phonons. The complete set of spectral densities for varying dimensions and coupling mechanism is shown in Appendix E.
@@
We now consider the model presented above to assess the performance of our approach on a more complex system. This model has a single controlled parameter $\bm{\theta}^{\mathrm{c}}=[\epsilon]$ which is varied such that each dot energy level is swept through the bias window. We choose the variable parameters to be $\bm{\theta}^{\mathrm{v}}=[\mu_{l},\mu_{r},\Gamma_{\mathrm{L}},\Gamma_{\mathrm{R}},t_{c},J_{0},T]$ which can be separated into differentiable parameters $[\Gamma_{\mathrm{L}},\Gamma_{\mathrm{R}},t_{c},J_{0},T]$ and non-differentiable parameters $[\epsilon_{0},\epsilon_{\mathrm{bias}}]$ which are the points where $\epsilon=0$ and $\epsilon=\mu_{\mathrm{L}}-\mu_{\mathrm{R}}$, respectively. The remaining parameters are set to representative fixed values of $c_{s}=3000\ \mathrm{ms}^{-1}$, $a=20\ \mathrm{nm}$, and $d=100\ \mathrm{nm}$. We choose these parameters to be fixed as they can all be estimated from the device design, but they could also be included in the set of differentiable parameters. The dot radius $a$ in particular has minimal impact on the fit for the energy scales probed by typical experimental measurements. The problem of estimating parameters for this model from a single current trace is expected to be underdetermined due to the number of parameters defining the system. We find fitting to data points is accurate, with an example shown in Figure 4. Discussion of parameter estimation for this model can be found in Appendix D.
Capitalising on the accuracy of fitting to data, we use our approach to characterise other aspects of a device hosting these dots. We focus on model selection for the interaction of the double quantum dot with phonons, using our fitting process to perform model selection on the phonon spectral density. The phonon spectral density is determined by material properties and dimensionality of the semiconductor in which the dots are defined, as well as the geometry of the dots, as outlined in Appendix E. Accurately distinguishing between piezoelectric and deformation potential phonon coupling would yield information about the device in which the dots are formed. Determining the dimensionality of the phonon spectral density would yield information about the confinement of phonon modes in the device and help identify any possible waveguide effects.
@@ VII Time Evolution
Having considered parameter estimation from steady-state solutions to quantum master equations using our differentiable solver, this section briefly explores time-evolution functionality and its limitations. Dynamic control of qubits is a natural motivation for studying time evolution, and thus we consider the model of a double quantum dot coupled to fermionic leads and a phonon bath outlined above as charge qubits have been implemented in such systems gorman2005charge ; petersson2010quantum . A Lindblad master equation can be recast as a set of coupled ordinary differential equations (ODEs) for time evolution solutions, which means standard ODE solvers can be used. Our solver implements a $4^{\mathrm{th}}$ order Runge-Kutta method butcher2016numerical to compute time-dynamics. In Appendix A, we compare performance of our ODE solver with direct computation of the steady state as used for previous results. To verify that our solver is accurate, Figure 7 shows that the steady state solution is reached after a suitable evolution time under the influence of a time-independent Hamiltonian. This figure also demonstrates batch computation of time-evolution, whereby a batch of $n_{\mathrm{b}}=250$ detuning values is evolved from an initial state of $|\psi(0)\rangle=(|0\rangle+|L\rangle+|R\rangle)/\sqrt{3}$ over 15 ns for 1500 time steps in 0.38s using the same hardware as discussed in Appendix A.
@@ VIII Conclusion
We have presented a fast and differentiable quantum master equation solver, and demonstrated its capability to compute steady state solutions and time evolution of open quantum systems. Our solver can be applied to any system described by a Lindblad master equation, and minimal changes would permit application to a more general time-local approach involving fewer approximations. The speed of our solver comes from its capability to compute a batch of outputs in a single evaluation of a model. This speed in turn facilitates more involved algorithms requiring many evaluations, which would be impractical using standard solvers. An example of such an algorithm is the characterisation approach we develop which employs both gradient-based and gradient-free optimisation, and MCMC methods.
Our characterisation algorithm has been applied to case studies of transport through a single quantum dot with an orbital excited state, and a double quantum dot coupled to a phonon bath. We have found our approach to be successful for dealing with parameter estimation in the case of a single quantum dot. To highlight the utility of our algorithm, we have also performed model selection to determine the electron-phonon coupling mechanism in a double quantum dot. Our approach is successful in reliably determining the ground truth spectral density. Our approach could be directly applied to experimental studies of the spectral density, such as in Ref. hofmann2020phonon . Decoherence from electron-phonon interactions impacts both charge and spin qubits, therefore characterising the details of this coupling mechanism is important for understanding qubit dynamics.
Our differentiable master equation solver and its application to parameter estimation contributes to the field of characterising quantum systems using differentiable models krastanov2020unboxing ; genois2021quantum ; youssry2024experimental . Our approach differs from existing methods by utilising significantly fewer data points and by dealing with a combination of differentiable and non-differentiable parameters. These attributes are key for the characterisation of more general systems where experimental data may be time-consuming or difficult to acquire. Physics-aware machine learning is a notable area in which differentiable models have been applied previously craig2024bridging ; pachalieva2022physics ; wright2022deep , and our solver stands to widen the impact of such approaches, particularly in the field of quantum transport.